# Kinetic theory of decentralized learning for smart active matter

### v1

@@ Abstract
Smart active matter describes agents which can process information according to their individual policy to solve predefined tasks. We introduce a theoretical framework to study a decentralized learning process in which agents can locally exchange their policy to optimize their reward and thus attain target macroscopic states. We use our formalism to derive explicit hydrodynamic equations for the policy dynamics. We apply the theory to two different microscopic models where policies correspond either to fixed parameters similar to evolutionary dynamics, or to state-dependent controllers known from the field of robotics. We find good agreement between theoretical predictions and agent-based simulations. By deriving fundamental control parameters and uncertainty relations, our work lays the foundations for a statistical physics analysis of decentralized learning.
@@
Decentralized optimization processes have been intensively studied in interdisciplinary fields using methods inspired from statistical physics, including the study of evolutionary and population dynamics in biology [1, 2, 3, 4] and the investigation of social processes [5], such as opinion dynamics [6]. In addition, decentralized adaptation raises a growing interest in robotics and engineering [7, 8, 9, 10, 11], in particular to find more stable policies than typical centralized protocols [9, 10]. Robustness and flexibility may be enhanced by drawing inspiration from collective animal behavior [12]. Although standard experiments on robotic swarms are in practice often limited to less than a hundred robots, numerical simulations may account for much larger swarms, calling for a statistical physics approach to describe assemblies of small robots interacting through information exchange [10].
@@
The idea of experimentally realizing large swarms of microrobots [13, 14, 15] as a way to build responsive or programmable metamaterials [16, 17, 18, 19] is also at the core of the emerging field of smart active matter [20, 21, 22, 23, 24]. Along this soft robotics perspective, one tries to build large numbers of extremely simplified microrobots [25, 26, 27, 28], whose size ranges from the granular scale [29, 30] down to the colloidal one [31, 19], and whose main features result from the physical properties of the soft material they are made of [32, 33, 34, 35]. These soft microrobots may integrate simple computation capabilities [36], or may be controlled by an external computer through machine learning using, e.g., reinforcement learning techniques [37, 38, 39, 40, 41, 42, 43, 44, 45, 46]. Although their starting points differ, both hard and soft robotics aim at building large assemblies of autonomous microrobots with communication and computation capabilities leading to adaptative collective behaviors [10, 11, 22, 23, 26, 47, 48]. This goal is also shared by the emerging topic of biological metamaterials, consisting for instance of insects aggregates [49, 50].
@@
The ongoing experimental and numerical development of smart active matter calls for appropriate theoretical approaches to describe its collective properties. Early contributions in this direction include models of individual [51, 52, 53, 54, 55] or collective navigation of active particles [56, 57], or of pedestrians in a dense crowd [58, 59]. Stochastic thermodynamic may also be relevant to describe smart active matter [60, 61]. Yet, apart from some first steps [60], a theoretical framework to describe the collective behavior of adaptive agents exchanging information with their neighbors, as found in experiments with small robots [7, 8, 9, 10, 11], is still lacking.
@@
In this Letter, we develop a statistical physics framework based on kinetic theory to describe decentralized learning in smart active matter with local information exchange. In contrast to usual active matter [62, 63, 64], smart active matter is made of agents having the ability to communicate, to respond to external stimuli and to adapt their behavior to attain targeted collective properties. The behavior of individual agents is determined by a ‘policy’, i.e., a set of rules controlling their dynamics, whose parameters can be tuned. Among the modeling approaches used to characterize active matter, kinetic theories have played an important role by providing a framework to derive macroscopic hydrodynamic equations from microscopic collision rules [65, 66, 67, 68, 69]. Here, we describe decentralized learning via the exchange of policies between neighboring agents, akin to collision rules. Such binary exchange rules set the stage for a kinetic theory framework, within which we systematically derive macroscopic hydrodynamic equations for the time-evolution of policies starting from agent-based models.
@@
Agent-based model of decentralized learning – We consider a generic system of $N$ agents in $d$ dimensions representing, e.g., small robots [10]. Agents are located at position $\bm{x}$, move at velocity $\bm{v}$ with an overdamped dynamics, and diffuse with a positional diffusion coefficient $D_{0}$. Agents sense their environment (e.g., external fields or properties of neighboring particles) to control their dynamical rules. The set of adjustable parameters controlling the dynamics, which are specific to each agent, is here generically called the agent’s policy $\mathcal{P}$. In the examples below, the policy either corresponds to the agent’s angular diffusion coefficient, or to the response of the self-propulsion speed to an external stimulus. The state $\mathcal{S}$ of an agent gathers internal physical variables (e.g., orientation, self-propulsion velocity, etc.) and sensors outputs such as measured light intensity. Agents can then process observed states and extract the signal $G(\mathcal{S})$, which fluctuates with a characteristic frequency $\tau_{G}^{-1}$. The signal can be stored in the ‘memory’ $\mathcal{M}$ via the running average $\mathcal{M}(t)$ through
$\frac{\text{d}\mathcal{M}}{\text{d}t}=-\frac{1}{\tau_{\mathcal{M}}}\,\big{(}\mathcal{M}-G(\mathcal{S})\big{)},$ (1)
on the memory time scale $\tau_{\mathcal{M}}$. All quantities $\mathcal{S}$, $\mathcal{M}$, and $\mathcal{P}$ may be multi-component vectors. To lighten the presentation of the general formalism, we mostly stick to the scalar case, but the multi-component vector formalism is presented in the supplemental material (SM) [70].
To enable learning, agents can exchange information about the efficiency of their individual policies to reach a targeted collective behavior. This efficiency is quantified by a reward function $\mathcal{R}(\mathcal{M})$, which rewards agents following the targeted collective behavior. Comparing its own reward with other agents enables each agent to learn improved policies (see Fig. 1 for an illustration). The overall goal of each agent is to learn and optimize a policy $\mathcal{P}$ maximizing the reward function. Agents are therefore randomly selected with a rate $\lambda_{E}=\tau_{E}^{-1}$, called the exchange rate ($\tau_{E}$ is the exchange time). Once selected, an agent $i$ iterates over all neighbors $n$, chosen in random order within the interactions radius $r_{c}$. For each selected neighbor $n$, agent $i$ becomes a teacher (and agent $n$ the student) with a probability $p_{T}(\mathcal{R}_{i},\mathcal{R}_{n})$ given by
@@
where $\mathcal{R}_{i}=\mathcal{R}(\mathcal{M}_{i})$ and $\mathcal{R}_{n}=\mathcal{R}(\mathcal{M}_{n})$, and $\alpha_{T}$ is a parameter. Otherwise, agent $n$ is the teacher and agent $i$ is the student. In either case, the student updates its policy $\mathcal{P}_{S}$ and memory $\mathcal{M}_{S}$ by taking the corresponding values of the teacher, $\mathcal{P}_{S}:=\mathcal{P}_{T}$ and $\mathcal{M}_{S}:=\mathcal{M}_{T}$. The memory of the student is updated together with the policy to make its reward consistent with the updated policy [10]. The parameter $\alpha_{T}$ in Eq. (2) controls the sharpness of $p_{T}$, and may be interpreted as the precision with which agents evaluate and communicate the reward $\mathcal{R}$. To enhance adaptability, we also include a small rate of spontaneous policy changes. By analogy with evolutionary processes in biology, we call these spontaneous changes ‘mutations’ [71] and model them as a diffusion process in policy space with diffusion coefficient $D_{M}$.
@@
Kinetic theory of decentralized learning – To obtain a statistical description of the above agent-based model and derive the time-evolution for the statistics of policies $\mathcal{P}$, we introduce a kinetic theory framework. Learning events between two agents are treated by analogy with binary collisions in usual kinetic theory. In this framework, an assembly of a large number $N$ of agents is described by a single-agent phase-space density $f(\mathcal{S},\mathcal{M},\mathcal{P},\bm{x},t)$ characterizing the probability to find an agent at time $t$ at position $\bm{x}$, in state $\mathcal{S}$, with memory $\mathcal{M}$ and policy $\mathcal{P}$. The time-evolution of $f(\mathcal{S},\mathcal{M},\mathcal{P},\bm{x},t)$ formally reads
$\frac{\partial f}{\partial t}=\mathcal{I}_{\rm phys}[f]+\mathcal{I}_{\rm sens}[f]+\mathcal{I}_{\rm mem}[f]+\mathcal{I}_{\rm pol}[f].$ (3)
@@
The term $\mathcal{I}_{\rm phys}[f]$ describes the physical part of agents dynamics, e.g., their motion, angular dynamics and possibly physical interactions, and $\mathcal{I}_{\rm sens}[f]$ defines the dynamics of the sensor output. The term $\mathcal{I}_{\rm mem}[f]$ in Eq. (3) encodes the memory dynamics given in Eq. (1). Finally, the term $\mathcal{I}_{\rm pol}[f]$ accounts for the dynamics of the policy $\mathcal{P}$, which includes diffusive mutations and communication-driven decentralized learning,
$\mathcal{I}_{\rm pol}[f]=D_{M}\frac{\partial^{2}f}{\partial\mathcal{P}^{2}}+\mathcal{I}_{\rm learn}[f].$ (4)
@@
The learning term $\mathcal{I}_{\rm learn}[f]$ represents the exchange of information between neighboring agents regarding their policies and the corresponding rewards. Mathematically, the term $\mathcal{I}_{\rm learn}[f]$ is a bilinear integral term accounting for collision-like, binary learning events. The explicit, but lengthy, expression of $\mathcal{I}_{\rm learn}[f]$ is reported in the End Matter (EM) section. We assume in this manuscript a time scale separation $\tau_{G}\ll\tau_{\mathcal{M}}\ll\tau_{E}$, i.e., a hierarchy of time scales separating signal dynamics $\tau_{G}$, memory $\tau_{\mathcal{M}}$ and exchange rate $\tau_{E}$. As a result, information advection by particle motion dominates over information propagation due to the finite communication range between agents. Within this approximation, our theory is thus purely local, however, it can be generalized to take into account the non-locality of communications or overlapping time scales.
@@
To characterize the policy statistics, we assume that the reduced phase-space density
@@
is approximately Gaussian and can be described by its lowest-order moments. We also define the policy-dependent average memory,
@@
The reduction steps from $f(\mathcal{S},\mathcal{M},\mathcal{P})$ to $\varphi_{0}(\mathcal{P})$ and $\mu_{\mathcal{M}}(\mathcal{P})$ are described in the EM section. To keep the problem tractable, we expand $\mu_{\mathcal{M}}(\mathcal{P},\bm{x},t)$ around a predefined policy value $\mathcal{P}_{E}$, which should be a rough estimate of the optimal policy, yielding (dropping $\bm{x}$ and $t$)
$\mu_{\mathcal{M}}(\mathcal{P})=\mu_{\mathcal{M}}^{(0)}+\mu_{\mathcal{M}}^{(1)}(\mathcal{P}-\mathcal{P}_{E})+\dots$ (7)
@@
For spatially uniform fields $\rho(t)$, $\mu(t)$ and $\sigma^{2}(t)$, the evolution equations take the form,
@@
$\frac{d\sigma^{2}}{dt}$ $=2D_{M}-\rho\sigma^{4}F_{2}\left(\mu,\sigma^{2};\mu_{\mathcal{M}}^{(n)}\right),$ (11)
@@
where $F_{m}\big{(}\mu,\sigma^{2};\mu_{\mathcal{M}}^{(n)}\big{)}$ are model-specific polynomials [70] depending on the parameters $\mu_{\mathcal{M}}^{(n)}(t)$ introduced in Eq. (7), which can be determined independently. The closed equations (10), (11) for the policy dynamics are one of the major results of this work. The full hydrodynamic equations including space-dependent fields are derived for two specific models in the SM [70]. Importantly, we can already draw some non-trivial general conclusions from Eqs. (10) and (11): (i) The derivative $\dot{\mu}(t)$ is proportional to the density $\rho(t)$ highlighting the many-body character of policy dynamics. Similarly, it is proportional to the diversity $\sigma^{2}(t)$, thus vanishing density or diversity imply vanishing adaptation rate $\dot{\mu}(t)$, consistent with Fisher’s fundamental theorem of natural selection [72]. (ii) On time scales $t>\tau_{\mathcal{M}}$ with convex fitness function, we have $F_{2}(\mu(t),\sigma^{2}(t);\mu_{\mathcal{M}}^{(n)}(t))>0$ and thus $\sigma^{2}(t)$ decays monotonically in the absence of mutations. (iii) This emphasizes the importance of mutations $D_{M}$ to maintain a sufficient level of diversity and thus enable efficient optimization. Qualitatively, mutations therefore play a role similar to driving in athermal physical systems, as they break detailed balance and maintain a steady-state dynamics at large time.
@@
Figure 2: Adaptation of the microswimmer model to a constant target velocity $V_{T}=0.85$: Results of kinetic theory (KT) and agent-based simulations (AS) without mutations ($D_{M}=0$) and with mutations ($D_{M}=0.1$). The vertical arrow indicates $\tau_{\mathcal{M}}=1$. (a) Average policy (rotation diffusion) $\mu(t)$; (b) diversity $\sigma^{2}(t)$. The black dashed-dotted curve is Eq. (12) fitted to the AS data with $D_{M}=0.1$.
@@
Phototactic, diffusive agents – As a first application of the above framework, we study a model of phototactic microswimmers ($d=2$) with decentralized learning capabilities. Each swimmer moves at speed $v_{0}$ in a direction given by its orientation angle $\theta$ in the $(x,y)$-plane, which defines the state $\mathcal{S}=\theta$. As for active Brownian particles [73], orientation can change randomly via rotational diffusion with strength $D_{\theta},$ but also via tumbling events towards the $x-$direction (direction of the light source) [74]. Physical interactions between microswimmers are neglected. The target state is a macroscopic flow along the $x$-axis in which the ensemble-averaged velocity $\bar{V}_{x}(D_{\theta})=\langle V_{x}\rangle_{D_{\theta}}$ is equal to $V_{T}<v_{0}$. This target state is achieved via a decentralized learning procedure in which each agent measures its velocity $G(\mathcal{S})=V_{x}=v_{0}\cos\theta$ and adapts its rotational diffusion coefficient $\mathcal{P}=D_{\theta}$ (i.e., its policy) by maximizing the reward function, $\mathcal{R}(\mathcal{M})=1-(\mathcal{M}-V_{T})^{2}$. Importantly, although physical interactions are neglected, learning leads to collective effects. The asymptotic solutions of Eqs. (10) and (11) (detailed derivations in [70]) read,
$\sigma^{2}(t)$ $=\sqrt{\frac{2D_{M}}{\lambda_{0}}}\frac{1}{\tanh(\sqrt{2D_{M}\lambda_{0}}(\tau_{0}+t))}$ (12)
$\mu(t)$ $=D_{\theta,T}-\frac{\mu_{0}\sqrt{2D_{M}\lambda_{0}}}{\sinh(\sqrt{2D_{M}\lambda_{0}}(\tau_{0}+t))},$ (13)

### v2

@@ Abstract
Smart active matter has the ability to control its motion guided by individual policies to achieve collective goals. We introduce a theoretical framework to study a decentralized learning process in which agents can locally exchange policies to adapt their behavior and maximize a predefined reward function. We use our formalism to derive explicit hydrodynamic equations for the policy dynamics. We apply the theory to two different microscopic models where policies correspond either to fixed parameters similar to evolutionary dynamics, or to state-dependent controllers known from the field of robotics. We find good agreement between theoretical predictions and agent-based simulations. By deriving fundamental control parameters and uncertainty relations, our work lays the foundations for a statistical physics analysis of decentralized learning.
@@
Decentralized optimization processes have been intensively studied in interdisciplinary fields using methods inspired from statistical physics, including the study of evolutionary and population dynamics in biology [1, 2, 3, 4] and the investigation of social processes [5], such as opinion dynamics [6]. In addition, decentralized adaptation raises a growing interest in robotics and engineering [7, 8, 9, 10, 11, 12], in particular to find more stable policies than typical centralized protocols [9, 11]. Robustness and flexibility may be enhanced by drawing inspiration from collective animal behavior [13]. Standard experiments on robotic swarms typically involve around a hundred robots, and their numbers are increasing. This growth calls for a statistical physics approach to describe assemblies of small robots interacting through information exchange [11].
@@
The idea of experimentally realizing large swarms of microrobots [14, 15, 16] as a way to build responsive or programmable metamaterials [17, 18, 19, 20] is also at the core of the emerging field of smart active matter [21, 22, 23, 24, 25, 26, 27]. Along this soft robotics perspective, one tries to build large numbers of extremely simplified microrobots [28, 29, 30, 31], whose size ranges from the granular scale [32, 33] down to the colloidal one [34, 20], and whose main features result from the physical properties of the soft material they are made of [35, 36, 37, 38]. These soft microrobots may integrate simple computation capabilities [39], or may be controlled by an external computer through machine learning using, e.g., reinforcement learning techniques [40, 41, 42, 43, 44, 45, 46, 47, 48, 49]. Although their starting points differ, both hard and soft robotics aim at building large assemblies of autonomous microrobots with communication and computation capabilities leading to adaptative collective behaviors [11, 12, 24, 25, 29, 50, 51]. This goal is also shared by the emerging topic of biological metamaterials, consisting for instance of insects aggregates [52, 53].
@@
The ongoing experimental and numerical development of smart active matter necessitates appropriate theoretical approaches to describe its collective properties. Early contributions in this direction include models of individual [54, 55, 56, 57, 58] or collective navigation of active particles [59, 60], or of pedestrians in a dense crowd [61, 62]. Stochastic thermodynamics may also be relevant to describe smart active matter [63, 64]. Yet, apart from some first steps [65, 63], a theoretical framework to describe the collective behavior of adaptive agents exchanging information with their neighbors [12, 11] is still lacking.
@@
In this Letter, we develop a kinetic theory framework to describe decentralized learning in smart active matter with local information exchange. In contrast to usual active matter [66, 67, 68], smart active matter consists of agents capable of responding to external stimuli, communicating and adapting their behavior. Through this adaptability, they can learn to solve specific tasks or to attain targeted collective properties. The behavior of individual agents is determined by a ‘policy’, i.e., a set of rules controlling their dynamics, whose parameters can be tuned. Among the modeling approaches used to characterize active matter, kinetic theories have played an important role by providing a framework to derive macroscopic hydrodynamic equations from microscopic collision rules [69, 70, 71, 72, 73]. We describe decentralized learning via the exchange of policies between neighboring agents, akin to collision rules, setting the stage for a kinetic theory framework, within which we systematically derive macroscopic hydrodynamic equations for the time-evolution of policies starting from agent-based models.
@@
Agent-based model of decentralized learning – We consider a generic system of $N$ agents in $d$ spatial dimensions representing, e.g., small robots [11]. Agents are located at position $\bm{x}$, move at velocity $\bm{v}$ with an overdamped dynamics, and diffuse with a positional diffusion coefficient $D_{0}$. Agents sense their environment (e.g., external fields or properties of neighboring particles) to control their dynamical rules. The set of adjustable parameters controlling the dynamics, whose values are specific to each agent, is here generically called the agent’s policy $\mathcal{P}$. In the examples below, the policy either corresponds to the agent’s angular diffusion coefficient, or to the response of the self-propulsion speed to an external stimulus. The state $\mathcal{S}$ of an agent gathers internal physical variables (e.g., orientation, self-propulsion velocity, etc.) and sensors outputs such as measured light intensity. Agents can then analyze observed states and extract the processed information $G(\mathcal{S})$, which fluctuates on a characteristic time scale $\tau_{G}$. The processed information is then filtered and stored in the ‘memory’ $\mathcal{M}$ via the running average on the time scale $\tau_{\mathcal{M}}$,
$\frac{\text{d}\mathcal{M}(t)}{\text{d}t}=-\frac{1}{\tau_{\mathcal{M}}}\,\big{(}\mathcal{M}(t)-G(\mathcal{S}(t))\big{)},$ (1)
where $t$ is time. The purpose of the memory is to extract the relevant signal from potentially noisy observations $G(\mathcal{S})$. This ensures that reward is not awarded for attaining one specific state but rather to promote long-time behavior [11]. Eq. (1) is equivalent to a convolution of the processed information $G(\mathcal{S})$ by an exponential memory kernel of characteristic time scale $\tau_{\mathcal{M}}$. The quantities $\mathcal{S}$, $\mathcal{M}$, and $\mathcal{P}$ may be multi-component vectors. To lighten the presentation of the formalism, we stick to the scalar case, but the multi-dimensional formalism is presented in the supplemental material (SM) [74].
To enable learning, agents can exchange information about the efficiency of their individual policies. This efficiency is quantified by a scalar reward function $\mathcal{R}(\mathcal{M}),$ which is evaluated solely based on each agent’s individual memory $\mathcal{M}$. Agents will thus optimize their personal reward, but the reward function may be chosen to promote a specific target collective behavior (see examples below). Comparing its own reward with other agents enables each agent to learn improved policies (Fig. 1). Agents are therefore randomly selected with rate $\lambda_{T}=\tau_{T}^{-1}$, called the teaching rate ($\tau_{T}$ is the teaching time). Once selected, an agent $i$ iterates over all neighbors $n$, chosen in random order within the interaction radius $r_{c}$. For each selected neighbor $n$, agent $i$ becomes a teacher (and agent $n$ the student) with a probability $p_{T}(\mathcal{R}_{i},\mathcal{R}_{n})$ given by
@@
where $\mathcal{R}_{i}=\mathcal{R}(\mathcal{M}_{i})$ and $\mathcal{R}_{n}=\mathcal{R}(\mathcal{M}_{n})$. Otherwise, agent $n$ is the teacher and agent $i$ is the student. In either case, the student updates its policy $\mathcal{P}_{S}$ and memory $\mathcal{M}_{S}$ by adapting the corresponding values of the teacher, i.e., $\mathcal{P}^{\text{new}}_{S}=\mathcal{P}_{T}$ and $\mathcal{M}^{\text{new}}_{S}=\mathcal{M}_{T}$. The memory of the student is updated together with the policy to make its reward consistent with the updated policy [11]. The parameter $\alpha_{T}$ in Eq. (2) controls the sharpness of $p_{T}$, and may be interpreted as the precision with which agents evaluate and communicate the reward $\mathcal{R}$. In Eq. (2) only reward differences matter and agents are lacking a concept of an absolute reward, which highlights the decentralized nature of the optimization. To enhance adaptability, we also include a small rate of spontaneous policy changes. By analogy with evolutionary processes in biology, we call these spontaneous changes ‘mutations’ [75] and model them as a diffusion process in policy space with diffusion coefficient $D_{\rm mut}$.
@@
Kinetic theory of decentralized learning – To obtain a statistical description of the above agent-based model and derive the time-evolution for the statistics of policies, we introduce a kinetic theory framework. This approach is natural since the learning events correspond to binary, intermittent interactions, which can be treated in analogy with binary collisions in usual kinetic theories. In this framework, an assembly of a large number $N$ of agents is described by a single-agent phase-space density $f(\mathcal{S},\mathcal{M},\mathcal{P},\bm{x},t)$ characterizing the probability to find an agent at time $t$ at position $\bm{x}$, in state $\mathcal{S}$, with memory $\mathcal{M}$ and policy $\mathcal{P}$. The time-evolution formally reads
$\frac{\partial f}{\partial t}=\mathcal{I}_{\rm phys}[f]+\mathcal{I}_{\rm mem}[f]+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{I}_{\rm learn}[f]}.$ (3)
@@
The term $\mathcal{I}_{\rm phys}[f]$ describes the physical dynamics of the state $\mathcal{S}$ and the position $\bm{x}$, e.g., advection and diffusion, angular dynamics and physical interactions. $\mathcal{I}_{\rm mem}[f]$ in Eq. (3) encodes the memory dynamics given in Eq. (1). Finally, the term $\mathcal{I}_{\rm learn}[f]$ accounts for the dynamics of the policy $\mathcal{P}$, which includes diffusive mutations and communication-driven decentralized learning,
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{I}_{\rm learn}[f]}=D_{\rm mut}\frac{\partial^{2}f}{\partial\mathcal{P}^{2}}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{I}_{\rm teach}[f]}.$ (4)
@@
The teaching term $\mathcal{I}_{\rm teach}[f]$ describes the binary teaching events introduced in the agent-based model. Mathematically, the term $\mathcal{I}_{\rm teach}[f]$ is a bilinear integral of the phase-space density $f$. The explicit, but lengthy, expression of $\mathcal{I}_{\rm teach}[f]$ is reported in the End Matter (EM) section. We assume in this manuscript a time scale separation $\tau_{G}\ll\tau_{\mathcal{M}}\ll\tau_{T}$, i.e., a hierarchy of time scales separating processed information dynamics $\tau_{G}$, memory $\tau_{\mathcal{M}}$ and teaching events $\tau_{T}$. In other words, agents have a memory time $\tau_{\mathcal{M}}$ long enough to filter the signal from noisy observations and are learning sufficiently slow to enable the memory to converge towards its steady-state value for a fixed policy. Information advection by particle motion therefore dominates over information propagation due to the finite communication range between agents. Within this approximation, our theory is thus purely local. However, it can be generalized to consider non-local communication or overlapping time scales.
@@
To characterize the policy statistics, we assume that $f$ is approximately Gaussian in $\mathcal{M}$ and $\mathcal{P}$, but this assumption can also be relaxed (see EM). Consequently, the reduced phase-space density
@@
can be described by its lowest-order moments. We also define the policy-dependent average memory,
@@
The reduction steps from $f(\mathcal{S},\mathcal{M},\mathcal{P},\bm{x},t)$ to $\varphi_{0}(\mathcal{P},\bm{x},t)$ and $\mu_{\mathcal{M}}(\mathcal{P},\bm{x},t)$ are described in the EM section. To derive closed hydrodynamic equations for the moments introduced above, we expand $\mu_{\mathcal{M}}(\mathcal{P},\bm{x},t)$ around a predefined policy value $\mathcal{P}_{*}$. To ensure that this approximation is consistent, $\mathcal{P}_{*}$ should be a rough estimate of the optimal policy, yielding
$\mu_{\mathcal{M}}(\mathcal{P},\bm{x},t)=\mu_{\mathcal{M}}^{(0)}(\bm{x},t)+\mu_{\mathcal{M}}^{(1)}(\bm{x},t)(\mathcal{P}-\mathcal{P}_{*})+\dots$ (7)
@@
For spatially uniform fields, $\rho(t)$, $\mu(t)$, and $\sigma^{2}(t)$, the evolution equations take the form,
@@
$\frac{d\sigma^{2}}{dt}$ $=2D_{\rm mut}-\rho\sigma^{4}F_{2}\left(\mu,\sigma^{2};\mu_{\mathcal{M}}^{(n)}\right),$ (11)
@@
where $F_{m}\big{(}\mu,\sigma^{2};\mu_{\mathcal{M}}^{(n)}\big{)}$ with $m=1,2$ are model-specific polynomials [74] depending on the parameters $\mu_{\mathcal{M}}^{(n)}(t)$ introduced in Eq. (7), which can be determined independently. The closed Eqs. (10) and (11) for the policy dynamics are one of the major results of this work. The full hydrodynamic equations including space-dependent fields are derived for two specific models in the SM [74]. An explicit model where space-dependence is taken into account is given below. Importantly, we can already draw some non-trivial general conclusions from Eqs. (10) and (11): (i) The time derivative $\dot{\mu}(t)$ is proportional to the density $\rho(t)$, highlighting the many-body character of policy dynamics. Similarly, it is proportional to the diversity $\sigma^{2}(t)$, thus vanishing density or diversity imply vanishing adaptation rate $\dot{\mu}(t)$, consistently with Fisher’s fundamental theorem of natural selection [76]. (ii) On time scales $t>\tau_{\mathcal{M}}$ with a convex fitness function, we have $F_{2}(\mu(t),\sigma^{2}(t);\mu_{\mathcal{M}}^{(n)}(t))>0$ and thus $\sigma^{2}(t)$ decays monotonically in the absence of mutations. (iii) This emphasizes the importance of mutations $D_{\rm mut}$ to maintain a sufficient level of diversity and thus enable efficient optimization. Qualitatively, mutations therefore play a role similar to driving in athermal physical systems, as they break detailed balance and maintain a steady-state dynamics at large time.
@@
Figure 2: Adaptation of the microswimmer model to a constant target velocity $V_{T}=0.85$: Results of kinetic theory (KT) and agent-based simulations (AS) without mutations ($D_{\rm mut}=0$) and with mutations ($D_{\rm mut}=0.1$). The vertical arrow indicates $\tau_{\mathcal{M}}=1$. (a) Average policy (rotation diffusion) $\mu(t)$. (b) Diversity $\sigma^{2}(t)$ (variance of policy). The black dashed-dotted curve is Eq. (12) fitted to the AS data with $D_{\rm mut}=0.1$.
@@
Phototactic, diffusive agents – As a first application of the above framework, we study a model of phototactic microswimmers (in $d=2$) with decentralized learning capabilities. Each swimmer moves at speed $v_{0}$ in a direction given by its orientation angle $\theta$ in the $(x,y)$-plane, which defines the state $\mathcal{S}=\theta$. As for active Brownian particles [77], orientation can change randomly via rotational diffusion with strength $D_{\theta},$ but also via tumbling events towards the $x-$direction (direction of the light source) [78]. Physical interactions between microswimmers are neglected. The target behavior is a macroscopic flow along the $x$-axis in which the ensemble-averaged velocity $\langle V_{x}\rangle=N^{-1}\sum_{i=1}^{N}V_{x,i}$ is equal to the target velocity $V_{T}$. This behavior can be promoted via a decentralized learning procedure in which each agent measures its own velocity $G(\mathcal{S})=V_{x}=v_{0}\cos\theta$ and adapts its rotational diffusion coefficient $\mathcal{P}=D_{\theta}$ (i.e., its policy). By maximizing the reward function, $\mathcal{R}(\mathcal{M})=-(\mathcal{M}-V_{T})^{2}$, each agent’s time-averaged local velocity, $\mathcal{M},$ converges towards $V_{T}$ and thus induces the same ensemble-averaged velocity. Importantly, although physical interactions are neglected, learning is collective due to information exchange. The asymptotic solutions of Eqs. (10) and (11) (detailed derivations in [74]) read,
$\sigma^{2}(t)$ $=\sqrt{\frac{2D_{\rm mut}}{\lambda_{0}}}\frac{1}{\tanh(\sqrt{2D_{\rm mut}\lambda_{0}}(\tau_{0}+t))}$ (12)
$\mu(t)$ $=D_{\theta,T}{+}\frac{\eta_{0}\sqrt{2D_{\rm mut}\lambda_{0}}}{\sinh(\sqrt{2D_{\rm mut}\lambda_{0}}(\tau_{0}+t))},$ (13)